OpenStudy (anonymous):

one leg of a right triangle is 3 feet longer than the other leg. The length of the hypotenuse is 15ft. Find the length of the other two sides. x^2(x+3)^2=(15)^2

OpenStudy (anonymous):

i know that 3^2 is 9= and 15^2=225

OpenStudy (anonymous):

\[x ^{2} + (x+3)^{2} = 15^{2}\]

OpenStudy (anonymous):

you got that part right it looks like...

OpenStudy (anonymous):

now solving for x?

OpenStudy (anonymous):

having troubles with that?

OpenStudy (anonymous):

yes im stuck i believe im missing a step to square i did the 3 and 15

OpenStudy (anonymous):

hmm the whole (x+3) is squared... that means (x+3)(x+3) .... you have to foil it...

OpenStudy (anonymous):

or expand it, however your text calls it...

OpenStudy (anonymous):

remember how to do that?

OpenStudy (anonymous):

im trying

OpenStudy (anonymous):

k I'll check it when you're done

OpenStudy (anonymous):

is it 2x+9....

OpenStudy (anonymous):

x*x + x*3 + 3*x +3*3

OpenStudy (anonymous):

x^2+x^3+3x+9

OpenStudy (anonymous):

x*3 = 3*x I just put them in that order to help you to see how the expanding goes...

OpenStudy (anonymous):

so x^2 + 3x +3x +9 (or x^2 +6x +9)

OpenStudy (anonymous):

Is that clear? Have you worked with quadratic equations much?

OpenStudy (anonymous):

this was the first lesson my professor breezed through it

OpenStudy (anonymous):

do you see how I got x^2 +6x +9 from (x+3)(x+3) ?

OpenStudy (anonymous):

yes you got x^2 by mult x by x... than 3+3 last 3 times 3

OpenStudy (anonymous):



OpenStudy (anonymous):

x*x is x^2 3*x is 3x another 3*x is another 3x and 3*3 is 9

OpenStudy (anonymous):

so... I hope that's settled:) now we need to add the other x^2 we had in your equation x^2 +x^2 +6x + 9 =15^2 2x^2 +6x +9 =15^2

OpenStudy (anonymous):

yes that last line is where I left off and got stuck!

OpenStudy (anonymous):

2X^2+6X+9=225

OpenStudy (anonymous):

yes, now we have to solve for x so set it equal to zero and either factor it or use the quadratic formula..

OpenStudy (anonymous):

turns out it's factorable...

OpenStudy (anonymous):

(x+12)(x ___) can you determine what goes in the blank?

OpenStudy (anonymous):

what times 12 multiplies to the product of 225 right???

OpenStudy (anonymous):

no, we set it equal to zero... 2x^2 + 6x +9 -225 = 225 -225 = 0 so it's 2x^2 +6x -216 =0 divide through by two: x^2 +6x -108 =0 then factor: (x+12)(x ___) =0

OpenStudy (anonymous):

Sorry, I typed 'divide through by zero' I meant "divide through by two" of course:)

OpenStudy (anonymous):

(x-9)

OpenStudy (anonymous):

good:)

OpenStudy (anonymous):

so (x+12)(x-9) =0 x= ....?

OpenStudy (anonymous):

-12 and...?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

-12 and 9

OpenStudy (anonymous):

yes. -12 is obviously not a useful answer here... so x=9!

OpenStudy (anonymous):

it's not useful because there cannot be any negative numbers?!

OpenStudy (anonymous):

yes. it's the side for a triangle we can't draw. x=9 is the side for a triangle we can draw:)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sure! be sure to put the length of both sides in your answer one side is 9 and the other side is 9+3 ... (you can verify that 9^2 +12^2 = 15^2 in order to check your work)

OpenStudy (anonymous):

okay i did they said the answer was correct

OpenStudy (anonymous):

great!